OpenStudy (anonymous):

the front row of the stadium has 25 seats. each of the other rows has 2 more seats than the row in front of it. how many seats are there altogether in the first 20 rows?

OpenStudy (goformit100):

Use Sum Formula of Arithmetic progression @neer2890

OpenStudy (anonymous):

what is the formula?

OpenStudy (goformit100):

Sir first you try

OpenStudy (anonymous):

its\[s _{n}=\frac{ n(a _{1}+a _{2}) }{ 2 }\]

OpenStudy (neer2890):

so in 1st row , 25 seats are there. in 2nd row, 25+2=27 seats are there. in 3rd row, 29 seats are there. .. .. .. for 20th row use formula \[a _{n}=a+(n-1)d\] where an is nth term a= 1st term d is common difference between any two terms

OpenStudy (anonymous):

but its asking how many there are altogether in the first 20 rows

OpenStudy (neer2890):

so\[a _{20}= 25+(20-1)2=25+19*2=63\] now the question becomes 25+27+29+.............+63 so use \[S _{n}=\frac{ n }{ 2 }[a+l]\] where, Sn is sum of n terms a is 1st term and l is last term.

OpenStudy (anonymous):

ok tten what about this problem: the first term of a sequence is 1, and the 5th term is 47. if the difference between consecutive terms is constant, what is the 2nd term?

OpenStudy (neer2890):

use the same formula \[a _{n}=a+(n-1)d\]

OpenStudy (anonymous):

what would be d though?

OpenStudy (neer2890):

\[a _{5}=a+(5-1)d\] find d and use \[a _{2}=a+(2-1)d\] to find a2

OpenStudy (anonymous):

but how do i find d?

OpenStudy (neer2890):

you know a5=47 and a=1 just put their values in a5=a+(5-1)d and find d.

OpenStudy (anonymous):

\[S _{n}=\frac{ n }{ 2 }\left\{ 2a+\left( n-1 \right) d \right\}\]

OpenStudy (anonymous):

a=25 d=2 n=20